### 08 Dec 2023 v1

# VISAGE: Video Instance Segmentation with Appearance-Guided Enhancement

Abstract
In recent years, online Video Instance Segmentation (VIS) methods have shown remarkable advancement with their powerful query-based detectors. Utilizing the output queries of the detector at the frame level, these methods achieve high accuracy on challenging benchmarks. However, we observe the heavy reliance of these methods on the location information that leads to incorrect matching when positional cues are insufficient for resolving ambiguities. Addressing this issue, we present  VISAGE that enhances instance association by explicitly leveraging appearance information. Our method involves a generation of queries that embed appearances from backbone feature maps, which in turn get used in our suggested simple tracker for robust associations. Finally, enabling accurate matching in complex scenarios by resolving the issue of over-reliance on location information, we achieve competitive performance on multiple VIS benchmarks. For instance, on YTVIS19 and YTVIS21, our method achieves 54.5 AP and 50.8 AP. Furthermore, to highlight appearance-awareness not fully addressed by existing benchmarks, we generate a synthetic dataset where our method outperforms others significantly by leveraging the appearance cue. Code will be made available at https://github.com/KimHanjung/VISAGE.
Figure 1: Qualitative results across challenging scenarios. Predicted results using query-propagation [16, 37], query-matching [17, 36, 29], and our appearance-guided methods. The first row illustrates a shot change across consecutive frames, a scenario where previous methods fail to maintain consistent tracking. The second and third rows demonstrate trajectory intersections, leading to id-switching with previous methods. Unlike previous methods, our method successfully tracks objects without switching or losses. Best viewed in color.
1 Introduction
Video Instance Segmentation (VIS) is a challenging task that requires classification, segmentation, and tracking of distinct instances throughout a video sequence [31]. Current studies in VIS can be primarily categorized into two approaches: online and offline, based on whether a video is processed in a per-frame or per-clip manner. Recently, the advancement of frame-level object detectors has resulted in online methods becoming increasingly dominant in the VIS field.
As detectors directly impact the accuracy in the video domain, recent online models are primarily built using the powerful query-based detectors [38, 5]. Spatially decoding image information, the detectors are designed to represent object-wise information using the queries. Therefore, the online VIS methods reuse these queries from the detectors and have achieved substantial improvements in multiple challenging benchmarks [25, 33] by mostly adopting either propagation [16, 23] or matching [17, 29, 36] strategies. However, tracking under complex scenarios such as shot changes or trajectory intersections (Fig. 1) are yet imperfect, resulting in the degradation of the overall accuracy.
Examining these failure cases, we observe that object-wise information of the queries are significantly imbalanced: heavy reliance on positional cues, and less reflection of appearances. As demonstrated in Fig. 1, previous query-based methods [16, 36, 17, 29, 37], display a tendency to maintain the spatial order of previous predictions in their current predictions. To support this argument, we conduct additional analyses by horizontally flipping images to generate two-frame pseudo videos. The existing models manifest association errors despite the distinct exterior patterns of objects (Fig. 2 Top), which highlight the dependence on object locations. As there exist multiple scenarios that cannot be fully handled with the imbalanced information, such a phenomenon necessitates the models to take object appearances into consideration.
We introduce VISAGE (Video Instance Segmentation with Appearance-Guided Enhancement), a method that leverages appearance cues as a crucial indicator for distinguishing instances. First, we present a lightweight branch for generating appearance queries that encapsulate the visual characteristics of an instance. Specifically, given mask predictions of object queries from the detector, this branch explicitly embeds appearance by performing mask pooling operations to backbone feature maps. Then, our method separately uses the contrastive loss [26, 3] to pairs of appearance queries and object queries, which improve the association of instances across a video by alleviating the position-appearance imbalance.
Additionally, we present a simple tracker that employs appearance cues as an additional indicator and leverages temporal information. Previous methods [29, 36] employ multiple refinement steps for removing redundant mask proposals and adopt handcrafted threshold values to improve the accuracy. While these intricate tracker modules raise the tracking quality, they become dependent on multiple hyperparameters that are heuristically optimized for each dataset. In order to alleviate such dependence, our method streamlines the tracker and dramatically reduces the number of hyperparameters, eliminating thresholding and instead incorporating appearance cues. Nonetheless, a lack of temporal information still exists, inherent in query-matching online methods, as they are only aware of adjacent frames. We address this limitation by using a simple memory bank to facilitate temporal awareness.
Our approach VISAGE has many desirable properties. Enhanced by the appearance guidance, VISAGE achieves robust tracking in complex scenarios, as illustrated at Fig. 1 (a) and bottom of Fig. 2. Furthermore, with its simplified tracker that effectively utilizes the past history of both object and appearance queries, VISAGE demonstrates competitive performance across all benchmark datasets. We validate our methodology through extensive evaluations on the four standard VIS benchmarks: YouTubeVIS-19/21/22[31, 32, 33] and OVIS [25]. Unfortunately, these datasets are mostly composed of videos with general scenarios, thus not suitable for precise analyses of the effectiveness of leveraging appearance embeddings. Therefore, to validate our method’s effectiveness in leveraging appearance cues, we develop a large-scale pseudo dataset and conduct comparisons of its performance with other methods. Finally, VISAGE demonstrates superior performance over previous methods [16, 36] in the comparison.
Figure 2: Proof of concept demonstrated with a flipped image. Previous methods [17, 36, 37, 16, 29] struggle with instance matching in flipped images, showing a dependency on location. Our method, VISAGE, addresses this by emphasizing appearance, enabling accurate instance matching even with image flipping.
2 Related Works
2.1 Online Video Instance Segmentation
Contrary to traditional online approaches [31, 34, 11, 19], modern online methods utilize query-based detectors with an emphasis on query association strategy. Query-based online methods can be divided into two main strategies: query-matching and query-propagation.
Query-matching approaches focus on dynamically building tracklets in an online manner. For instance, MinVIS [17] utilizes Mask2Former [5] as its query-based detector trained exclusively at the image level. During inference, MinVIS independently predicts outputs at the frame level and then associates them through bipartite matching between frame queries. To enhance the distinctiveness of queries, IDOL [29] introduces contrastive learning, inspired by QDTrack [8]. They also adopt same inference process with QDTrack, additionally propose temporal weighted softmax with memory bank for leveraging temporal information. However, those memory bank is not used at training stage. CTVIS [36] further improves upon IDOL by removing the discrepancy between training and inference caused by the use of the memory bank. By utilizing the memory bank during training, CTVIS eliminates this discrepancy and incorporates historical instances. All these methods rely on the outputs of query-based detectors for their matching processes. In contrast, our approach VISAGE, introduces additional query to ensure more stable matching.
Figure 3: Overview of  VISAGE. (a) The proposed  VISAGE’s architecture which generate object embedding and appearance embedding. (b) Overall inference pipeline of  VISAGE: At time step $t-1$, the memory bank is updated with both the appearance embedding and the object embedding. Then, at time step $t$, the memory embedding is read from the memory bank and used for matching. (c). Details of the matching process: In that scenario, using only object embeddings leads to incorrect matching. However, when guided by the appearance embedding, the matching process can be corrected. Best viewed in color.
Query-propagation methods utilize output queries from previous frames to detect objects in the current frame. InsPro [12] propagates their instance query and proposal for association and proposes duplicate removal strategy. CAROQ [6] extends Mask2Former [5] for application in VIS methods and propagates its output queries along with relative object queries. GenVIS [16] employs an offline VIS model [15], and adapts it to a query-propagation approach, notably omitting heuristic association strategies. This method reduces computational overhead by freezing the backbone network and query-based detector, thus effectively handling longer video sequences. Similarly, DVIS [37] introduces referring cross-attention, which is designed to capture the correlation between the current and previous frames using their queries. Yet, these methods are limited by their local matching strategy, which focuses on aligning tracklets with ground truths. This approach often leads to unstable tracking outcomes, as it does not consider the full video context. TCOVIS [23] addresses this issue by shifting from the local matching strategy of GenVIS to a global matching approach, thereby achieving more robust tracking across videos.
2.2 Offline Video Instance Segmentation
Previous offline VIS methods [1, 30] treat video clip input as a spatio-temporal volume. However, with the advent of query-based detectors, recent offline methods have shifted to treating videos through queries. Recent offline methods mostly leverage query-based detectors. VisTR [27] builds upon DETR [2] to create an end-to-end framework. However, the self-attention mechanism in VisTR is computationally heavy. To address this, IFC [18] introduces memory tokens that allow communication between frames, reducing the computational overhead. SeqFormer [28] restructures the decoder to process each frame separately, enhancing the model’s sensitivity to instance movement by capturing each location. Mask2Former-VIS [4] extends the Mask2Former [5] for video-level tasks, achieving notable performance improvements. TeViT [35] presents a novel backbone that improves temporal information processing, leveraging the capabilities of ViT [7]. VITA [15] effectively aggregates video information by learning the relationships between the frame-level queries.
3 Method
We propose a novel online VIS framework that leverages appearance information to reduce the heavy reliance on location information. Our newly introduced appearance-aware query guides the association procedure, ensuring correct matching. Finally, we introduce a simple inference method that aligns with our design principles, without heuristics.
3.1 Query-based Detector
In most recent query-based object detection frameworks, the architecture is typically segmented into three integral components: the backbone, the transformer encoder, and the transformer decoder. Initially, the backbone generates the low-level image feature maps, which encapsulate the essential visual information. Subsequent to this, the transformer encoder employs self-attention mechanisms to enhance the feature representation, as detailed in [2, 38]. Finally, in the transformer decoder, the identified objects are decoded into $N$ learnable queries.
Our method uses the well-established query-based detection framework [5], maintaining the original structure with the backbone, transformer encoder, and transformer decoder. Furthermore, we utilize the feature maps from the backbone to create the appearance query, and we use the output of the transformer decoder as the object query.
3.2 Appearance-Guided Enhancement
Before the emergence of query-based tracking approaches, video tracking methods [31, 8, 20] generate instance features from backbone feature maps. This is achieved using operations like RoIPool [10] and RoIAlign [14], followed by associating them across the video. In line with these traditional methods, we use average pooling to extract appearance feature from the backbone feature maps, guided by the predicted masks as shown in Fig. 3 (a). Therefore, in addition to the object queries already in use, we further employ the appearance query as an additional indicator. We transform both appearance queries and object queries, leading to the creation of appearance embeddings $\mathbf{e}_{a}\in\mathbb{R}^{N\times C}$ and object embeddings $\mathbf{e}_{i}\in\mathbb{R}^{N\times C}$, respectively.
Our appearance embeddings enhance the matching process when using only object embeddings leads to incorrect matches, as illustrated in Fig. 3 (c). Relying solely on the similarity of object embeddings may lead to ambiguity due to their similar positions across subsequent frames. However, this ambiguity can be resolved by also considering the similarity of appearance embeddings, as their distinct appearances provide additional discriminative information. With this guidance, we leverage both appearance embeddings $\mathbf{e}_{a}$ and object embeddings $\mathbf{e}_{i}$ to identify the optimal match between queries across subsequent frames as shown in Algorithm 1 (lines 4-10).
To ensure more distinctive embeddings for better association, we apply contrastive loss to both object and appearance embeddings, similar to the approach used in previous methods [29, 8]. We sample key frame and reference frame, then calculate the loss for each set of embeddings separately. Consequently, our final loss is a weighted summation of original query-based detector’s loss and the contrastive loss for both the appearance embeddings and object embeddings.
Algorithm 1 Inference pipeline of VISAGE.
1:model $\mathcal{F}$, frames $\{x^{t}\}_{t=1}^{T}$
2:predictions $\mathcal{P}$
3:$\mathcal{P}\leftarrow\{\}$
${}\mathcal{M}\leftarrow\{\}$
5:$\texttt{idx}\leftarrow[0,N-1]$
6:for $t$ in $[1,T]$:
7:     ${\mathbf{p}}^{t},\mathbf{e}_{i}^{t},\mathbf{e}_{a}^{t}\leftarrow\mathcal{F}(x^{t})$
8:     $\mathcal{M}\leftarrow\mathcal{M}+(\mathbf{e}_{i}^{t}[\texttt{idx}],\mathbf{e}_{a}^{t}[\texttt{idx}])$
[…]
10:     $\mathcal{P}\leftarrow\mathcal{P}+{\mathbf{p}}^{t}[\texttt{idx}]$
11:     $\mathbf{s}\leftarrow\texttt{cos_sim}(\mathbf{e}_{i}^{t},\mathbf{m}_{i}^{t})+\texttt{cos_sim}(\mathbf{e}_{a}^{t},\mathbf{m}_{a}^{t})$
[…]
13:end for
14:return $\mathcal{P}$
3.3 Inference with Appearance
As shown in Algorithm 1, our tracker employs a simple yet effective matching process that aligns the current object embeddings and appearance embeddings with their respective counterparts from the previous frame. We compute the similarity scores for each object-appearance embedding pair using Cosine Similarity, following a method similar to that in [17]. Then, employing the sum of object and appearance similarities, we utilize the Hungarian algorithm [21] (linear_sum_assignment in Algorithm 1) to achieve optimal assignment.
In addition, we incorporate a simple memory bank to compensate for the lack of temporal information, a limitation inherent in online processes, as used in previous methods [29, 36]. We stack the states of previous queries from the most recent frames within our memory bank, which has a size of $W$. From this bank, we read a memory embedding $m\in\mathbb{R}^{N\times C}$ using the read_memory() function in Algorithm 1. Specifically, this function is implemented by temporally weighting the embeddings to put more emphasis on recent queries while utilizing the confidence scores for selective weighting. For the current memory embedding $m^{t}$, the calculation of the weighting for each embedding at the previous time step $w\in[1,W]$ can be formally expressed as follows:
$\mathbf{m}^{t}=\sum\limits_{w=1}^{W}\left(\mathbf{e}^{t-w}s^{t-w}\times\frac{W}{w}\right),$ (1)
where $s$ denotes the confidence score. This memory embedding represents the object and appearance, denoted as $m^{t}_{i}$ and $m^{t}_{a}$ respectively, in Algorithm 1.
Finally, we sequentially associate the predictions ($\mathbf{p}$ in Algorithm 1) from each frame using the obtained assignment. By employing a simple inference pipeline coupled with an efficient memory bank, we introduce an effective approach.
4 Experiments
Method Setting YouTube-VIS 2019 YouTube-VIS 2021
[…]
IFC [18] offline 41.2 65.1 44.6 42.3 49.6 35.2 55.9 37.7 32.6 42.9
Mask2Former-VIS [4] offline 46.4 68.0 50.0 - - 40.6 60.9 41.8 - -
MinVIS [17] online 47.4 69.0 52.1 45.7 55.7 44.2 66.0 48.1 39.2 51.7
IDOL [29] online 49.5 74.0 52.9 47.7 58.7 43.9 68.0 49.6 38.0 50.9
VITA [15] offline 49.8 72.6 54.5 49.4 61.0 45.7 67.4 49.5 40.9 53.6
GenVIS [16] online 50.0 71.5 54.6 49.5 59.7 47.1 67.5 51.5 41.6 54.7
GenVIS [16] offline 51.3 72.0 57.8 49.5 60.0 46.3 67.0 50.2 40.6 53.2
DVIS [37] online 51.2 73.8 57.1 47.2 59.3 46.4 68.4 49.6 39.7 53.5
DVIS [37] offline 52.6 76.5 58.2 47.4 60.4 47.4 71.0 51.6 39.9 55.2
TCOVIS [23] online 52.3 73.5 57.6 49.8 60.2 49.5 71.2 53.8 41.3 55.9
CTVIS [36] online 55.1 78.2 59.1 51.9 63.2 50.1 73.7 54.7 41.8 59.5
VISAGE online 54.5 77.2 60.3 50.9 61.7 50.8 72.9 55.6 43.3 58.8
Table 1: Comparisons on the YouTube-VIS 2019 and 2021 validation sets. Methods are denoted as online or offline, indicated by the text color. We highlight the best performance in bold and the second-best performance with an underline.
4.1 Datasets
We evaluate VISAGE on four VIS benchmarks: Youtube-VIS (YTVIS) 2019/2021/2022 [31, 32, 33] and Occluded VIS (OVIS) [25]. The YTVIS datasets contain 40 predefined categories in their videos. YTVIS 2019 is the first and largest dataset for video instance segmentation. It includes 2,238 videos for training, 302 for validation, and 343 for the testing. YTVIS 2021 expands upon the YTVIS 2019 dataset, containing 2,985 videos for training, 421 for validation, and 453 for testing, while refining annotations and modifying some categories. Additionally, to represent challenging scenarios with longer sequences, YTVIS 2022 is added. It adds 71 extra videos to the validation set of YTVIS 2021.
OVIS stands out for its more complex and longer videos, featuring 25 categories and comprising 2,238 videos for training, 302 for validation, and 343 for testing. In comparison to YTVIS, it contains a higher number of instances per video, averaging 5.8, and a total of 296k masks. Additionally, the average length of videos in OVIS is approximately 12 seconds.
4.2 Implementation Details
We adopt the Mask2Former [5] as our query-based detector. All of our models are initialized with parameters pre-trained on COCO dataset [24] with ResNet-50 [13] backbone. We also adopt COCO joint training following previous works [28, 15, 16, 36, 23]. Our batch includes 16 videos. In our experimental setup, we set the weights for the contrastive losses applied to both the appearance and object embeddings at $2.0$. We follow our detector’s loss weight configurations as detailed in  [5] for other losses in our setup. The window size of memory bank $W$ is set to 5 during inference. Finally, our method is trained using 4 NVIDIA A6000 GPUs.
[…]
MinVIS [17] 23.3 47.9 19.3 20.2 28.0
VITA [15] 32.6 53.9 39.3 30.3 42.6
GenVIS [16] 37.5 61.6 41.5 32.6 42.2
TCOVIS [23] 38.6 59.4 41.6 32.8 46.7
VISAGE 36.7 60.0 41.1 35.4 45.1
Table 2: Comparison on the YTVIS 2022 long videos set. Bold and underline highlight the highest and second-highest performances, respectively. For methods that are both online and semi-online, the online version is used.
4.3 Main Results
We compare our methods with state-of-the-art methodsg on four VIS benchmarks: YTVIS 2019/2021/2022 and OVIS. The results are reported in Tab. 1, Tab. 2 and Tab. 3.
Youtube-VIS 2019 & 2021.
As shown in Tab. 1, we compare our VISAGE with state-of-the-art methods. On the YTVIS 2019 benchmark, VISAGE demonstrates performance comparable to the highest-performing method [36]. Notably, both our VISAGE method and the best-performing method [36], surpass previous methods by a large margin. In the YTVIS 2021 benchmark, an improved version of YTVIS 2019, VISAGE outperforms other existing methods. We achieve competitive performance on both the YTVIS 2019 and 2021 datasets, by incorporating appearance-guided enhancement.
Youtube-VIS 2022.
The results for the YTVIS 2022 benchmark are presented in Tab. 2. This benchmark contains longer videos compared to previous benchmarks (YTVIS 2019/21). Our VISAGE exhibits slightly lower performance compared to GenVIS [16] and TCOVIS [23]. These methods utilize the entire history of previous information by propagating the queries. Nevertheless, our method employs only a fixed number of previous frames and shows comparable performance.
[…]
VITA [15] 19.6 41.2 17.4 11.7 26.0
MinVIS [17] 25.0 45.5 24.0 13.9 29.7
IDOL [29] 30.2 51.3 30.0 15.0 37.5
GenVIS [16] 35.8 60.8 36.2 16.3 39.6
DVIS [37] 31.0 54.8 31.9 15.2 37.6
TCOVIS [23] 35.3 60.7 36.6 15.7 39.5
CTVIS [36] 35.5 60.8 34.9 16.1 41.9
VISAGE 34.8 59.8 35.3 16.2 39.4
Table 3: Comparison on the OVIS validation set. Bold and underline indicate the best and second-best performances, respectively. We use the online version for GenVIS [16] and DVIS [37].
[…]
Transformer Encoder 52.8 73.8 58.1 49.4 60.2
Backbone 54.5 77.2 60.3 50.9 61.7
Table 4: Ablation study of the target feature maps which generate appearance query
OVIS.
Our method, VISAGE, is evaluated on the OVIS dataset. In Tab. 3, VISAGE demonstrates competitive performance when compared to other state-of-the-art methods [16, 23, 36]. It is important to note that these methods incorporate sequential information during training. Consequently, these methods shows marginally better performance than VISAGE, which is expected given that the OVIS dataset is characterized by its longer video sequences.
4.4 Ablation Studies
In this section, we provide the ablation studies for our proposed method and discuss its effect. All ablation experiments are conducted on YTVIS 2019 [31] validation set.
Appearance feature.
In Tab. 4, we analyze the impact of feature maps which create the appearance query. When we make the appearance query from the transformer encoder feature maps, there is a slight drop in performance. This indicates that backbone feature maps contain rich visual information compare to transformer encoder feature maps.
Appearance guidance.
Tab. 5 shows effectiveness of our appearance-guided enhancement. We omit appearance information by not computing the appearance similarity in line 9 of Algorithm 1, resulting in a matching process that relies solely on object similarity. As indicated by rows 1 and 3 in Tab. 5, the absence of appearance information leads to a degradation in performance. Further analysis of the effectiveness of appearance guidance is discussed in Sec. 4.5.
Memory bank.
In Tab. 5, we demonstrate the necessity of the memory bank for compensating for the lack of temporal information. We evaluate performance differences between scenarios with and without the use of the memory bank. Rows 1 and 2 in Tab. 5 show decreased performance compared to rows 3 and 4, respectively. If there is no memory bank, our method lacks access to historical information beyond the frame right before the current one, since it measures similarities solely between the current frame and the one immediately preceding it. By implementing a memory bank, our method gains awareness of previous frames, which leads to improved performance.
Appearance Memory Bank YouTube-VIS 2019
[…]
49.9 71.4 54.7 47.0 58.7
50.5 71.6 55.5 47.4 60.6
53.4 76.8 58.7 49.8 61.2
54.5 77.2 60.3 50.9 61.7
Table 5: Ablation study on appearance guidance and memory bank utilization in YTVIS 2019 validation Set, with memory bank window size $W=5$.
[…]
[29, 36] (w/o NMS) 52.7 75.3 57.6 49.3 61.6
[29, 36] (w/ NMS) 54.4 77.9 59.0 49.9 62.8
ours (w/o NMS) 54.5 77.2 60.3 50.9 61.7
Table 6: Ablation study of the tracker type on YTVIS 2019 validation set.
Simplified tracker.
Previous query-matching based methods [29, 36], extend the tracker from [8], which employs heuristic post-processing techniques such as NMS. In contrast, our tracker relies solely on a cosine similarity metric and the Hungarian algorithm for matching. To analyze the impact of such handcrafted design, we modify our inference pipeline to align with the approaches used in previous methods [29, 36]. As demonstrated in Tab. 6, both the previous strategies and our simplified version show comparable performance. However, removing the heuristic design element NMS from the previous approach results in performance degradation. This outcome implies that previous approaches rely on handcrafted design, unlike our simpler tracker.
4.5 Analysis of Appearance-Guided Enhancement
T-SNE Visualization.
As illustrated in Fig. 4, we analyze the impact of appearance-guided enhancement in VISAGE on its association capabilities using the OVIS dataset by visualizing the query embeddings for an entire video, with consistent colors indicating the same instance. The visualization includes three types of query embeddings: object-only, appearance-only, and a combination of both object and appearance queries. The embeddings that combine object and appearance query result in a well-clustered visualization. While the appearance-only query embeddings exhibit reasonable clustering, the integration of both types of queries leads to more distinct and pronounced clustering. This enhanced clustering indicates that our appearance-guided enhancement leads to better association.
Figure 4: T-SNE visualization on the OVIS dataset. Each row representing three different videos. Each column corresponds to the type of query embedding utilized. Points plotted in the same color indicate the same instance across the dataset. Best viewed in color.
Pseudo Dataset.
Traditional datasets do not fully cover the complex scenarios that our approach is designed to address. To validate VISAGE in a more intuitive manner, we construct a pseudo dataset consisting of synthetic videos. Instances in pseudo videos are created by compositing objects from the COCO dataset [24] with Copy-and-Paste augmentation [9]. The background images for the pseudo dataset are randomly sourced from BG-20k [22]. Additionally, we employ Bezier Curves to simulate the movement of objects. It includes two types of videos: track, where objects move following arbitrary bezier curves, and swap, where objects’ locations are randomly swapped along their trajectories. Except for the movement of instances, both types of datasets are generated under the same conditions.
As shown at the top of Fig. 5, instances in the track type pseudo video move along a Bezier curve. Consequently, complex scenarios, such as intersections between instances or movements out of the frame, occur naturally. On the other hand, the swap type presents a challenge for methods that primarily rely on location information, as this dependence results in incorrect matching. This is illustrated at the bottom of Fig. 5, where the positions of instances in the pseudo video are suddenly swapped. Such scenarios verify the method’s awareness of appearance cues when location information is no longer a reliable indicator.
Figure 5: Visualization of the pseudo dataset. In track type videos, instances move along random bezier curves. On the other hand, the swap type refers to a scenario where the positions of each instance are exchanged in the middle of the video. The colored dot above each instance represents the corresponding instance in the swapped frame.
[…]
Track GenVIS [16] 54.8 70.6 58.8 61.2 65.9
CTVIS [36] 63.7 78.7 69.7 70.3 74.8
VISAGE (w/o app) 65.2 80.9 71.3 70.2 75.2
VISAGE 65.8 81.1 71.7 70.8 76.3
Swap GenVIS [16] 41.8 59.7 44.0 50.8 57.0
CTVIS [36] 53.1 71.7 56.9 61.7 65.7
VISAGE (w/o app) 51.8 72.1 54.4 57.7 62.8
VISAGE 63.0 79.5 69.3 67.8 73.5
Table 7: Comparisons on Pseudo dataset. Bold denote the highest accuracy.
In  Tab. 7, we conduct an evaluation of various online VIS methods [16, 36] including VISAGE on the pseudo dataset using published ResNet-50 backbone weights, which are trained on the YouTube-VIS 2019 dataset. All of these methods use COCO joint training, ensuring a fair comparison. The overall trend is similar to that observed on the YTVIS 2019 dataset. In contrast, VISAGE outperforms other methods in swap type videos, mainly due to the challenges posed by the dataset in associating instances using only location information. Consequently, other methods demonstrate degraded performance on swap type videos compared to their performance on track types. Furthermore, the absence of appearance information in the matching process notably impacts VISAGE’s performance on swap type videos, although it remains relatively unaffected for track type videos. It proves that appearance cue serves as an additional indicator, providing robustness in scenarios with swapped object positions.
Figure 6: Qualitative results of VISAGE. Videos are sourced from OVIS [25] and YTVIS 2021 [32] datasets. These videos represent complex scenarios, characterized by intersections and reappearances. Best viewed in color.
4.6 Qualitative Results
Fig. 6 displays the visualization results of our VISAGE across various challenging scenarios. In the first row, several cats are positioned closely together and the red-colored cat crosses another cat. In such cases, relying solely on location information can often result in identity-switching errors. However, our VISAGE tracks each cat accurately despite their adjacent positions and sequential intersections. In the second row, we present a similar scenario involving different classes: a person and a cow.
In the third row, we showcase another challenging scenario where an object completely disappears and then reappears: a cat disappears and then reappears. Our memory bank, which stores the previous queries of the cat, enables us to reestablish the association when the cat reappears. The fourth row demonstrates a similar scenario, where our memory bank proves effective: a small sedan is completely occluded by the truck. Even though the large truck in the foreground and the small, disappeared car are located closely, their distinct appearances ensure accurate matching.
5 Limitations and Future Works
VISAGE addresses the insufficiency of existing query-based online VIS methods, specifically their lack of appearance awareness. However, our approach has two limitations. First, it assumes equal importance for appearance and location information for its simplicity. In some scenarios, appearance may be more critical than location, while in others, location may be prioritized. Designing a method to adaptively weight the importance of appearance and location information for stable tracking will be a future direction. Second, VISAGE focuses exclusively on the query-matching approach, and we observe that query-propagation methods also tend to rely on location information. Another future direction involves incorporating appearance information awareness into query-propagation approaches.
6 Conclusion
In this paper, we explore the importance of appearance in tracking objects, an aspect often taken for granted yet overlooked by current online VIS methods. We simply leverage the appearance cue when tracking objects, enabling more effective distinction of objects in challenging scenarios. Our method, while simple yet effective, achieves performance comparable to previous methods across various VIS benchmarks. However, as existing VIS benchmarks do not focus on appearance-requiring scenarios, we generate a synthetic dataset that necessitates the use of appearance information. This dataset serves to validate our appearance-aware approach and our approach surpasses other methods with very large margin. We believe that recognizing and leveraging the importance of appearance can lead to progress of VIS.